# All-weather Multi-Modality Image Fusion: Unified Framework and 100k Benchmark

### v2

@@ III-C1 PFB
where denotes the convolution operator, and represents the Hadamard product. Subsequently, by introducing the matrix-vector forms of , , , , and , which are denoted as K, J, I, A, and D, respectively, Eq. 5 can be redefined as follows:
@@
In adverse weather conditions like dense haze, rain, or snow, light undergoes scattering and absorption, causing alterations in both image reflectance and illumination. For example, haze reduces the intensity of light reflected from objects to the camera by scattering it, with this effect being more pronounced at greater distances. Rain and snow also alter the surface reflectivity of objects when they adhere to them, affecting how light is reflected. Additionally, atmospheric particles can cause light to be reflected, refracted, and absorbed, impacting the overall illumination of the scene. Consequently, for the prediction of the feature , by analyzing changes in reflectance and illumination of objects in an image, we can infer changes in light transmission, thus realizing estimation of the image transmission information.
We disassemble into a combination of reflectance and illumination features, which is modeled as follows:
where and are the reflectance and illumination filters of , respectively. Subsequently, we compute the reflectance and illumination feature responses by solving the following optimization problem:
@@
The goal of RFB is used to compute the reflectance feature. According to [19], we fix to update by rewriting Eq. 9 as:
@@
where is the stack of feature responses , is the update of at the -th iteration, and are the learnable convolutional layers related to the filters , and represents the soft thresholding operator. Similar to SFB, we expand the iterations in Eq. 11 into a network as shown in Fig. 2.
@@
The goal of IFB is to compute the illuminance feature in a hazy image, the structure of this network is similar to that of DFB. By fixing to update , Eq. 9 can be rewritten as:
@@
III-D Loss Function
The loss function integrates two components. The first part focuses on the loss function used to ensure the quality of the fusion results, while the second part addresses the loss function constraining the generation of reflectance and illumination features. The fusion loss includes structural similarity index (SSIM), mean square error (MSE) and color consistency loss. The calculation for SSIM can be expressed as follows:
@@ III-D Loss Function
Colour consistency loss is used to ensure that the fused image maintains consistent colours with the . We transform the image into the YCbCr color space and apply a constraint using the Euclidean distance between the Cb and Cr channels. This can be expressed as:
@@
The total loss of AWFusion is denoted as:
@@
III-E MMIF Benchmark for Adverse Weather
@@
Current MMIF datasets lack comprehensive scene coverage under adverse weather. Combining existing datasets (RoadScene[28], MSRS[46], M3FD[3], and LLVIP[47]) and our newly collected data in real-world scenes, we have established a benchmark AWMM-100k, which encompasses a diverse range of weather conditions. Each type of weather (rain, haze, and snow) is categorized into heavy, medium, and light levels. Specific descriptions are shown in Table I. The AWMM-100k dataset was collected using a DJI M30T drone, equipped with high-resolution visible and thermal imaging cameras to capture data in complex scenes.
@@
Collecting a large amount of aligned paired multi-modality data and corresponding ground truth in real adverse weather is challenging, leading most current research to rely on synthetic data for training. Effectively degrading clean data and imbuing the model with adaptability to real scenes are pivotal for constructing benchmark. However, the degradation of infrared images in severe weather remains a current research gap. To address this, we analyze data collected during real severe weather to determine the most effective methods for degrading infrared images. Observations of the data in Fig. 3 allow us to draw the following conclusions: 1) While the visible image exhibits degradation due to adverse weather conditions, the corresponding infrared image lacks explicit weather features, showcasing only varying degrees of degradation commensurate with the extent of image disturbance, such as diminished contrast and detail. 2) The depth of the scene affects the extent of light refraction and scattering, resulting in greater blurring in distant scenes compared to closer ones. Based on the above conclusions, we propose an adaptive degradation strategy for infrared images based on scene depth. First, we transpose the image into the Fourier domain and use the depth map as the weights for degradation. Subsequently, the weight map is normalised and multiplied with the amplitude and phase of the image respectively and then inverted to obtain the degraded image. However, distant pixels in the depth map are always represented in black, i.e., the pixel value is . In order to suppress too small weights, we add , and to points with pixel values less than according to the three degrees of light, medium and heavy, respectively.
@@
Degraded visible images generation process is illustrated in Fig. 1. We employ a phenomenological pipeline [48] to introduce haze into source images, while rain and snow images are generated using Photoshop software utilizing masks based approach [49]. In addition, Fig. 4 presents examples of multi-modality data captured in real-world scenarios under adverse weather. Moreover, this benchmark provides labels for object detection task. These attributes render it a versatile resource for investigating a broad spectrum of visual tasks.
IV Experiment
IV-A Experimental Configurations
@@ IV-B Training Details
For the training process, three specialized teacher networks and one student network were trained under different weather conditions. For different adverse weather conditions, we selected 1,000 pairs of multi-modality images from the AWMM-100k dataset as the training set for the teacher network. Each image was randomly cropped to a size of pixels to form a batch. The initial learning rate for the Adam optimizer was set to , and the model training was conducted over 300 iterations. During the knowledge transfer phase, the teacher network remained fixed while the student network was trained for 200 epochs on a mixed training set comprising three weather types. Furthermore, we employed two LLRR blocks, while the remaining parameters were set according to the original paper. The number of LCSC blocks iterated in the RFB, IFB, and SFB models is four, with 24 filters per block. The size of the convolutional kernel used in each convolutional layer is nine. All experiments were carried out using the PyTorch framework on a system equipped with two NVIDIA 4090 GPUs.
IV-C Qualitative Comparison
IV-C1 Comparison on synthetic data
@@
Figs. 5, 6 and 7 present the fusion results of different algorithms on the synthetic dataset. Notably, only the proposed algorithm performs fusion under adverse weather conditions, while the comparison methods are fused in ideal scenes. Additionally, we include pseudo-color images of the disturbed visible images and the fusion results from the student model to further demonstrate the effectiveness of the proposed algorithm in mitigating severe weather interference. AWFusion-T and AWFusion-S represent the teacher and student models of the proposed algorithm, respectively. As shown in Figs. 5, 6 and 7, all comparison methods achieve good fusion results in ideal scenes. Four methods—U2Fusion, LRRNet, MURF, and IGNet—produce fusion results with low contrast. In contrast, CoCoNet and MUFusion enhance the source image features, leading to some redundancy and loss of detail in the thermal radiation information in the fusion results. In daytime scenes, the TIM, TarDal, and DRMF methods effectively retain the texture information from the visible images. However, in nighttime scenes, the TarDal method exhibits a loss of detail. Additionally, both MUFusion and CrossFuse show poor color retention of the sky, introducing excessive infrared information into the region, which is not compatible with the human visual system.
@@
IV-C2 Comparison on real-world data
@@
Fig. 9, we present the fusion results of all methods in a real haze scene. Additionally, we evaluate the fusion performance of different methods using two downstream tasks: object detection and depth estimation. We used YOLOv7 [10] as a detector and MiDas [57] as a depth estimator. Unlike synthetic data, the input images used to comparison methods and the proposed algorithm are the same, as there is no corresponding ground truth for real scene data. Furthermore, we retrained three high-performing dehaze algorithms (Dehamer[58], DehazeFormer[59], UCL-Dehaze[60]) using images from the AWMM-100K to optimize the visible images for input into the fusion comparison methods. However, as seen in Fig. 8, all dehaze methods resulted in varying degrees of color distortion, suggesting that feeding such distorted inputs into the different fusion methods could lead to unfair comparisons. Therefore, we provided all fusion methods with uniform input images.
@@
As shown in Fig. 9, in the first set of fusion comparison tests, CoCoNet, TarDal, and MUFusion effectively retain the significant thermal radiation information from the infrared map. These three methods, which tend to enhance infrared data to achieve higher contrast, can address some challenges in real haze scenes. However, this approach excessively masks details from the visible image, preventing the recovery of clear scene details. Additionally, for pedestrians submerged in haze, DRMF and CrossFuse lose target information from the infrared image, incorrectly prioritizing the blurred pedestrian details from the visible image. In contrast, the proposed algorithm strikes an effective balance between the complementary information from different modalities, maintaining good overall scene contrast while also recovering some detail in unknown real scenes. By examining the accuracies of different methods in the object detection task, it is clear that only the AWFusion-S results successfully detect all target information in the scene. While the CoCoNet method enhances target information, this also introduces detection errors, suggesting that blind enhancement is not an effective solution to the problem of degraded fusion. The proposed algorithm effectively recovers scene details and integrates complementary multi-modality information, achieving the highest detection accuracy across all objects. For the depth estimation task, the results visualize each ability of method to extract overall scene information. While all methods preserved depth in the foreground, many, such as U2Fusion, LRRNet, CoCoNet, and MUFusion, struggled with accurate depth estimation for distant pedestrians obscured by haze. Additionally, haze affected the performance of TIM, IGNet, and DRMF, as these methods failed to maintain contrast between the roadside fence and its surroundings, leading to edge blurring in the depth map.
@@
In summary, the proposed algorithm outperforms current SOTA methods in both fusion performance and semantic information retention when addressing real-world scene challenges. It demonstrates excellent performance in downstream tasks, such as object detection and depth estimation, indicating its high practical application value in real-world.
IV-D Quantitative Comparison
IV-D1 Fusion Performance Comparison

### v3

@@ 3.3.1 Physical-Aware Clear Feature Prediction Block (PFB)
where denotes the convolution operator, and represents the Hadamard product. Subsequently, by introducing the matrix-vector forms of , , , , and , which are denoted as , , , , and , respectively, Eq. 10 can be redefined as follows:
@@
In adverse weather conditions like dense haze, rain, or snow, light undergoes scattering and absorption, causing alterations in both image reflectance and illumination. For example, haze reduces the intensity of light reflected from objects to the camera by scattering it, with this effect being more pronounced at greater distances. Rain and snow also alter the surface reflectivity of objects when they adhere to them, affecting how light is reflected. Additionally, atmospheric particles can cause light to be reflected, refracted, and absorbed, impacting the overall illumination of the scene. Consequently, for the prediction of the feature , by analyzing changes in reflectance and illumination of objects in an image, we can infer changes in light transmission, thus realizing estimation of the transmission information.
We disassemble into reflectance and illumination components, modeled as:
where and are the convolutional filters for reflectance and illumination, respectively.
To derive a tractable optimization formulation from the convolutional decomposition in Eq. 13, we first express the convolutions in matrix-vector form. Each filter and is converted into a Toeplitz matrix and , while the corresponding feature maps and are vectorized. This allows the decomposition to be equivalently written as a linear combination of reflectance and illumination components:
Then, we introduce reconstruction fidelity terms for both components, measured by the -norm of the difference between and their reconstructions:
Lastly, to encourage sparsity in the learned reflectance and illumination features and suppress redundant activations, we add an regularization term for each feature map. Combining the fidelity and sparsity terms leads to a joint optimization problem that can be solved using convolutional sparse coding.
Accordingly, the reflectance and illumination feature responses are obtained by solving:
Furthermore, We emphasize that our objective is adverse weather restoration rather than strict Retinex-based intrinsic decomposition. In our framework, reflectance and illumination serve as two feature subspaces for estimating transmission-related cues in the atmospheric scattering model. Although a shared sparsity prior is applied, convolutional filters and loss fusion still guide the model to learn distinct responses, effectively separating scene details from low-frequency degradation.
@@
The RFB is designed to extract reflectance features, we formalize this as a convolutional sparse coding problem. According to [10], we fix to update by rewriting Eq. 16 as:
@@
where is the stack of feature responses , is the update of at the -th iteration, and are the learnable convolutional layers related to the filters , and represents the soft thresholding operator. Similarly, we expand the iterations in Eq. 18 into a network as shown in Fig. 2.
@@
The goal of IFB is to compute the illuminance feature in the degraded image, the structure of this network is similar to that of RFB. By fixing to update , Eq. 16 can be rewritten as:
@@
3.4 Loss Function
The loss function consists of two main components: the first corresponds to the training loss of the teacher networks, while the second pertains to the training loss of the student network. The former is primarily based on standard image fusion losses, whereas the latter incorporates knowledge distillation.
3.4.1 Teacher Networks Loss Function
This loss function integrates two components. The first part focuses on the loss function used to ensure the quality of the fusion results, while the second part addresses the loss function constraining the generation of reflectance and illumination features. The fusion loss includes structural similarity index (SSIM) [59], mean square error (MSE) and color consistency loss. The calculation for SSIM can be expressed as follows:
@@ 3.4.1 Teacher Networks Loss Function
Colour consistency loss [68] is used to ensure that the fused image maintains consistent colours with the . We transform the image into the YCbCr color space and apply a constraint using the Euclidean distance between the Cb and Cr channels. This can be expressed as:
@@
The total loss of teacher network is denoted as:
3.4.2 Student Networks Loss Function
The student network learns from teacher networks trained for three types of adverse weather via knowledge distillation. Its training objective builds on the teacher network loss while incorporating a Collaborative Knowledge Transfer loss () and a Soft Contrastive Regularization loss ()[5].
The consists of the Projection Feature Error loss () and Projection Feature Validation losses . projects teacher and student feature maps into a shared lower-dimensional space to mitigate the impact of architectural differences on knowledge transfer, while ensures that the feature error computed in is based on informative and valid representations. In essence, this loss does not treat teacher features as hard targets; instead, the projection and validation mechanisms address feature space misalignment across networks. The treats outputs from well-performing teacher networks as positive (soft) targets and the original degraded images as negative samples, enabling the student network to learn effective weather-specific restoration solutions and facilitating smoother knowledge integration during early training.
In summary, the total loss of student network is denoted as:
@@ 3.4.2 Student Networks Loss Function
3.5 MMIF Benchmark for Adverse Weather
Dataset Overview
@@ 3.5 MMIF Benchmark for Adverse Weather
Current MMIF datasets lack comprehensive scene coverage under adverse weather. Combining existing datasets (RoadScene[63], MSRS[51], M3FD[34], and LLVIP[15]) and our newly collected data in real-world scenes, we have established a benchmark AWMM-100k, which encompasses a diverse range of weather conditions. Each type of weather (rain, haze, and snow) is categorized into heavy, medium, and light levels. Specific descriptions are shown in Table 1. The AWMM-100k dataset was collected using a DJI M30T drone, equipped with visible and thermal imaging cameras.
Challenges in Data Collection
@@
Collecting a large amount of aligned paired multi-modality data and corresponding ground truth in real adverse weather is challenging, leading most current research to rely on synthetic data for training. Effectively degrading clean data and imbuing the model with adaptability to real scenes are pivotal for constructing benchmark. However, the degradation of infrared images in severe weather remains a current research gap. To address this, we analyze data collected during real severe weather to determine the most effective methods for degrading infrared images. Observations of the data in Fig. 3 allow us to draw the following conclusions:
Visible images show clear weather-related degradation, while infrared images exhibit only general contrast and detail reduction without explicit weather cues.
The depth of the scene affects the extent of light refraction and scattering, resulting in greater blurring in distant scenes compared to closer ones.
Building on these observations, we proposed a adaptive degradation strategy for multi-modality images.
Infrared Degradation Strategy
First, we transpose the image into the Fourier domain and use the depth map as the weights for degradation. Subsequently, the weight map is normalised and multiplied with the amplitude and phase of the image respectively and then inverted to obtain the degraded image. However, distant pixels in the depth map are always represented in black, i.e., the pixel value is . In order to suppress too small weights, we add , and to points with pixel values less than according to the three degrees of light, medium and heavy, respectively.
Visible Image Degradation Strategy
@@
Degraded visible images generation process is illustrated in Fig. 1. We employ a phenomenological pipeline [61] to introduce haze into source images, while rain and snow images are generated using Photoshop software utilizing masks based approach [4]. To ensure the realism of synthetic rain and snow images in the AWMM-100K dataset, we extract masks from real-world rain and snow video sequences, which are then used in the rendering process to generate degraded images. This approach preserves the natural shape, orientation, motion blur, density, and spatial structure of raindrops and snowflakes, such as linear streaks for rain and scattering or aggregation patterns for snow [4, 26]. Consequently, the synthetic data exhibit high statistical and visual realism, enabling the model to generalize effectively to real-world scenarios. As shown in Fig. 4, the degradations in the synthetic data of AWMM-100K closely match those observed in real degraded scenes. Fig. 5 presents examples of multi-modality data captured in real-world scenarios. Moreover, this benchmark provides labels for object detection. These attributes make it a versatile resource for exploring a wide range of visual tasks.
4 Experiment
4.1 Experimental Setup
@@ 4.2 Training Details
For the training process, three specialized teacher networks and one student network were trained under different weather conditions. For different adverse weather conditions, we selected 1,000 pairs of multi-modality images from the AWMM-100k dataset as the training set for the teacher network. Each image was randomly cropped to a size of pixels to form a batch. The initial learning rate for the Adam optimizer was set to , and the model training was conducted over 300 iterations. During the knowledge transfer phase, the teacher network remained fixed while the student network was trained for 200 epochs on a mixed training set comprising three weather types. Furthermore, we employed two LLRR blocks, while the remaining parameters were set according to the original paper. The number of LCSC blocks iterated in the RFB, IFB, and SFB models is four, with 24 filters per block. The size of the convolutional kernel used in each convolutional layer is nine. All the experiments were conducted on a system equipped with two NVIDIA 4090 GPUs.
Source Method Pub. Background Car Person Bike Curve Car Stop Guardrail Color cone Bump mIoU U2Fusion[73] TPAMI 2020 98.39 92.88 73.25 73.03 52.82 73.75 86.2 40.31 83.2 74.87 LRRNet[23] TPAMI 2023 98.51 93.18 69.62 73.89 65.18 77.95 84.59 35.95 88.96 76.43 MURF[64] TPAMI 2023 99.13 95.56 83.54 86.65 76.27 88.21 93.48 76.25 90.59 87.74 TIM[39] TPAMI 2024 98.22 91.17 67.94 69.4 59.53 67.74 83.42 30.59 85.59 72.62 CoCoNet[35] IJCV 2023 98.3 92.03 69.31 70.66 57.72 73.62 87.49 29.44 84.9 73.72 TarDal[34] CVPR 2022 98.39 92.29 70 68.47 62.07 74.59 87.9 35.65 85.42 75.01 IGNet[25] MM 2023 99.12 95.61 83.09 87.22 74.97 87.96 93.1 76.13 90.9 87.6 DRMF[47] MM 2024 99.16 95.14 82.36 88.2 84.27 90.68 91.67 80.35 92.08 89.32 MUFusion[8] InfFus 2023 99.23 95.17 83.84 88.43 84.1 88.9 92.94 80.7 86.12 88.82 GT CrossFuse[22] InfFus 2024 98.04 90.3 67.43 68.51 58.58 71.81 68.13 29.92 68.7 69.04 Rain 99.32 96.49 84.77 88.76 86.35 90.45 93.91 78.45 91.82 90.04 Haze 99.19 95.69 81.51 81.94 85.21 89.5 91.57 72.73 93.57 87.88 Snow AWFusion-S - 99.25 95.92 81.24 86.4 86.77 90.36 93.74 78.04 92.32 89.34
Source Method Pub. People Car Bus Lamp Motorcycle Truck mAP@0.5 mAP@[0.5:0.95] U2Fusion[73] TPAMI 2020 0.816 0.903 0.901 0.716 0.711 0.801 0.808 0.512 LRRNet[23] TPAMI 2023 0.785 0.909 0.912 0.764 0.721 0.808 0.817 0.484 MURF[64] TPAMI 2023 0.815 0.885 0.881 0.711 0.675 0.8 0.794 0.498 TIM[39] TPAMI 2024 0.731 0.898 0.878 0.657 0.656 0.769 0.765 0.48 CoCoNet[35] IJCV 2023 0.791 0.889 0.889 0.696 0.641 0.762 0.778 0.49 TarDal[34] CVPR 2022 0.792 0.877 0.857 0.585 0.62 0.713 0.741 0.462 IGNet[25] MM 2023 0.793 0.859 0.836 0.532 0.554 0.721 0.716 0.439 DRMF[47] MM 2024 0.699 0.897 0.896 0.787 0.697 0.784 0.793 0.494 MUFusion[8] InfFus 2023 0.803 0.906 0.908 0.734 0.694 0.802 0.808 0.507 CrossFuse[22] InfFus 2024 0.692 0.891 0.871 0.671 0.674 0.771 0.762 0.473 GT AWFusion-S - 0.791 0.922 0.93 0.804 0.726 0.821 0.832 0.534
4.3 Qualitative Comparison
4.3.1 Comparison on synthetic data
@@
Figs. 6, 7 and 8 present the fusion results of different algorithms on the synthetic dataset. Notably, only the proposed algorithm performs fusion under adverse weather conditions, while the comparison methods are fused in ideal scenes. Additionally, we include pseudo-color images of the disturbed visible images and the fusion results from the student model to further demonstrate the effectiveness of the proposed algorithm in mitigating severe weather interference. AWFusion-T and AWFusion-S represent the teacher and student models of the proposed algorithm, respectively. As shown in Figs. 6, 7 and 8, all comparison methods achieve good fusion results in ideal scenes. Four methods, including U2Fusion, LRRNet, MURF, and IGNet, produce fusion results with low contrast. In contrast, CoCoNet and MUFusion enhance the source image features, leading to some redundancy and loss of detail in the thermal radiation information in the fusion results. In daytime scenes, the TIM, TarDal, and DRMF methods effectively retain the texture information from the visible images. However, in nighttime scenes, the TarDal method exhibits a loss of detail. Additionally, both MUFusion and CrossFuse show poor color retention of the sky, introducing excessive infrared information, which is not compatible with the human visual system.
@@
To demonstrate the capability of proposed method in multi-modal feature interaction and extraction, we conduct a comparison under degradation-free conditions. As shown in Fig. 9, all methods receive identical inputs. The results indicate that our method still maintains superior fusion performance and does not fail when applied to clean scenes, despite being trained on images affected by adverse weather.
4.3.2 Comparison on real-world data
@@
Fig. 11, we present the fusion results of all methods in a real haze scene. Additionally, we evaluate the fusion performance of different methods using two downstream tasks: object detection and depth estimation. We used YOLOv7 [55] as a detector and MiDas [43] as a depth estimator. Unlike synthetic data, the input images used to comparison methods and the proposed algorithm are the same, as there is no corresponding ground truth for real scene data. Furthermore, we retrained three high-performing dehaze algorithms (Dehamer[11], DehazeFormer[45], UCL-Dehaze[58]) using images from the AWMM-100K to optimize the visible images for input into the fusion comparison methods. However, as seen in Fig. 10, all dehaze methods resulted in varying degrees of color distortion, suggesting that feeding such distorted inputs into the different fusion methods could lead to unfair comparisons. Therefore, no restoration algorithm was incorporated into the fusion method to remove interfering factors from the source images.
@@
As shown in Fig. 11, in the first set of fusion comparison tests, CoCoNet, TarDal, and MUFusion effectively retain the significant thermal radiation information from the infrared map. These three methods, which tend to enhance infrared data to achieve higher contrast, can address some challenges in real haze scenes. However, this approach excessively masks details from the visible image, preventing the recovery of clear scene details. Additionally, for pedestrians submerged in haze, DRMF and CrossFuse lose target information from the infrared image, incorrectly prioritizing the blurred pedestrian details from the visible image. In contrast, the proposed algorithm strikes an effective balance between the complementary information from different modalities, maintaining good overall scene contrast while also recovering some detail in unknown real scenes. By examining the accuracies of different methods in the object detection task, it is clear that only the AWFusion-S results successfully detect all target information in the scene. While the CoCoNet method enhances target information, this also introduces detection errors, suggesting that blind enhancement is not an effective solution to the problem of degraded fusion. The proposed algorithm effectively recovers scene details and integrates complementary multi-modality information, achieving the highest detection accuracy across all objects. For the depth estimation task, the results visualize each ability of method to extract overall scene information. While all methods preserved depth in the foreground, many, such as U2Fusion, LRRNet, CoCoNet, and MUFusion, struggled with accurate depth estimation for distant pedestrians obscured by haze.
@@
In summary, the proposed algorithm outperforms current SOTA methods in both fusion performance and semantic information retention when addressing real-world scene challenges, and demonstrates excellent performance in downstream tasks. It demonstrates excellent performance in downstream tasks, such as object detection and depth estimation, indicating its high practical application value in real-world.
Methods Pub. Restor. ↑ ↑ ↑ ↑ ↓ ↑ ↑ ↑ ↑ Rain Text-DiFuse [69] NIPS 24 0.2442 0.3645 0.5283 0.3909 777.6507 0.1962 1.2938 0.3329 6.9609 ControlFusion [48] NIPS 25 0.3507 0.4750 0.6901 0.4361 739.1370 0.2996 1.1971 0.4576 7.0261 GIFNet [7] CVPR 25 0.2975 0.3686 0.8040 0.4834 672.4981 0.3474 1.3440 0.3544 5.5197 SAGE [60] CVPR 25 0.2286 0.5086 0.7878 0.3964 465.1593 0.2734 1.2986 0.3670 5.4524 OmniFuse [70] TPAMI 25 AdaIR [9] 0.2191 0.4027 0.5863 0.3774 549.3715 0.1771 1.0588 0.3340 6.9013 AWFusion-T 0.3580 0.6855 0.8444 0.4613 464.7867 0.3693 1.6311 0.4839 6.3160 AWFusion-S _ _ 0.3334 0.6428 0.8335 0.4691 476.3436 0.3528 1.6010 0.4618 6.2777 Haze Text-DiFuse [69] NIPS 24 0.3137 0.3463 0.6622 0.4391 597.2099 0.3035 1.2850 0.4021 7.3050 ControlFusion [48] NIPS 25 0.3841 0.4032 0.7309 0.4652 547.6077 0.3331 1.1490 0.4767 6.9131 GIFNet [7] CVPR 25 0.3124 0.3181 0.6878 0.4784 847.9477 0.3463 1.2118 0.3478 6.2780 SAGE [60] CVPR 25 0.3416 0.3948 0.7347 0.4553 557.8396 0.3316 1.1922 0.4336 6.6047 OmniFuse [70] TPAMI 25 AdaIR [9] 0.2638 0.3448 0.6483 0.4226 514.9563 0.2558 1.0287 0.3446 7.0612 AWFusion-T 0.4497 0.5889 0.7949 0.4838 374.9215 0.4078 1.2209 0.5585 6.9558 AWFusion-S _ _ 0.4283 0.5471 0.7738 0.4610 462.8585 0.3901 1.1719 0.5322 6.9469 Snow Text-DiFuse [69] NIPS 24 0.3047 0.3633 0.6263 0.4264 502.4444 0.2821 1.5002 0.4130 7.0752 ControlFusion [48] NIPS 25 0.4184 0.4610 0.7487 0.4855 427.3205 0.3554 1.3418 0.5194 7.0112 GIFNet [7] CVPR 25 0.3282 0.3305 0.7706 0.4854 482.2309 0.3722 1.3885 0.3882 6.0665 SAGE [60] CVPR 25 0.3149 0.4752 0.7910 0.4421 275.8248 0.3474 1.4294 0.4585 6.0712 OmniFuse [70] TPAMI 25 AdaIR [9] 0.2729 0.3845 0.6254 0.4114 318.5334 0.2515 1.1877 0.3977 7.0768 AWFusion-T 0.3580 0.5591 0.8179 0.4827 246.8077 0.3780 1.5700 0.4985 6.6173 AWFusion-S _ _ 0.3552 0.5535 0.8119 0.4798 270.1595 0.3734 1.4937 0.4937 6.6531
Methods Pub. Restor. ↑ ↑ ↑ ↑ ↓ ↑ ↑ ↑ ↑ Rain Text-DiFuse [69] NIPS 24 0.2485 0.3622 0.5280 0.4076 770.4863 0.2000 1.2776 0.3384 6.9881 ControlFusion [48] NIPS 25 0.3809 0.4677 0.6769 0.4420 740.8956 0.2956 1.1693 0.4531 7.0698 GIFNet [7] CVPR 25 0.3075 0.3747 0.8062 0.4804 670.4288 0.3512 1.3463 0.3650 5.5507 SAGE [60] CVPR 25 0.2355 0.5356 0.7925 0.4072 465.1666 0.2872 1.3023 0.3793 5.5068 OmniFuse [70] TPAMI 25 MODEM [57] 0.2248 0.4068 0.5903 0.3912 545.6923 0.1845 1.0391 0.3434 6.9319 AWFusion-T 0.3580 0.6855 0.8444 0.4613 464.7867 0.3693 1.6311 0.4839 6.3160 AWFusion-S _ _ 0.3334 0.6428 0.8335 0.4691 476.3436 0.3528 1.6010 0.4618 6.2777 Haze Text-DiFuse [69] NIPS 24 0.3513 0.3752 0.7024 0.4728 516.7434 0.3380 1.3969 0.4577 7.3805 ControlFusion [48] NIPS 25 0.4402 0.4761 0.7560 0.4981 553.4258 0.3673 1.2045 0.5511 7.1785 GIFNet [7] CVPR 25 0.3530 0.3366 0.7344 0.5007 1090.9966 0.3834 1.2959 0.4085 6.3176 SAGE [60] CVPR 25 0.4026 0.5124 0.7906 0.4826 626.5782 0.3935 1.3033 0.5216 6.6250 OmniFuse [70] TPAMI 25 MODEM [57] 0.2866 0.3705 0.6982 0.4621 529.6896 0.2834 1.0791 0.4028 7.1742 AWFusion-T 0.4497 0.5889 0.7949 0.4838 374.9215 0.4078 1.2209 0.5585 6.9558 AWFusion-S _ _ 0.4283 0.5471 0.7738 0.4610 462.8585 0.3901 1.1719 0.5322 6.9469 Snow Text-DiFuse [69] NIPS 24 0.3089 0.3672 0.6289 0.4312 492.1872 0.2848 1.4871 0.4207 7.0754 ControlFusion [48] NIPS 25 0.4150 0.4619 0.7486 0.4816 422.5371 0.3538 1.3377 0.5220 7.0211 GIFNet [7] CVPR 25 0.3227 0.3350 0.7693 0.4821 487.3009 0.3665 1.3761 0.3874 6.0588 SAGE [60] CVPR 25 0.3172 0.4913 0.7937 0.4453 262.5657 0.3486 1.4287 0.4653 6.0729 OmniFuse [70] TPAMI 25 MODEM [57] 0.2763 0.3902 0.6289 0.4182 311.3442 0.2540 1.1914 0.4052 7.0759 AWFusion-T 0.3580 0.5591 0.8179 0.4827 246.8077 0.3780 1.5700 0.4985 6.6173 AWFusion-S _ _ 0.3552 0.5535 0.8119 0.4798 270.1595 0.3734 1.4937 0.4937 6.6531
4.4 Quantitative Comparison
4.4.1 Fusion Performance Comparison